OpenStudy (anonymous):

Jim has grades of 84, 65, and 76 on three math tests. What grade must he obtain on the next test to have an average of exactly 80 for the four tests? A. 87 B. 95 C. 85 D. 92

OpenStudy (anonymous):

so if x is the grade he must obtain, then: (84+65+76+x)/4=80 from there (84+65+76+x)=320 now solve for x

OpenStudy (campbell_st):

well let x be the score in the next test so and he must score an average of 80 \[\frac{ 84 + 65 + 76 + x}{4} = 80\] to solve the 1st step is multiply both sides of the equation by 4 hope this helps